# Topological phases in discrete stochastic systems

### v1

@@ III.3 Formalism and mapping to quantum systems
Topological invariants in stochastic systems can be identified using results developed in quantum systems. In Ref. 35, a topological index was identified from the winding number in Fourier space of the phase associated to the determinant of the tilted current matrix associated to the bulk. In follow-up work, this index was further related to the topological index of the Hermitian Hamiltonian of an associated supersymmetric quantum system [34].
A more direct method that flexibly maps many systems described by the master equation to a quantum Hamiltonian was used in Ref. 33, although the resulting Hamiltonian in this case is typically non-Hermitian [71, 72]. Namely, the similarity of the master equation to the Schrödinger equation is exploited, as they are formally identical up to the constant . In quantum systems, topological invariants such as the Chern number are typically calculated from the eigenvectors of the Hamiltonian . When mapping a master equation to a corresponding Schrödinger equation, the eigenvectors of the transition matrix will be identical to those of , in the case of periodic boundary conditions. Hence, the same calculation that verifies if the quantum system is topological using its eigenvectors, can be performed for from a stochastic system. Using the bulk-boundary correspondence, a topological invariant that characterizes the bulk will predict a boundary response for the system. Further works have shown that this bulk-boundary correspondence has additional constraints in stochastic systems as compared to quantum ones [73, 74, 75], which is discussed below in Section VI.
@@ V.1.1 Long timescales: the circadian rhythm
A detailed case study of a topological mechanism in a concrete biological system is that of the KaiABC system for the circadian clock of cyanobacteria [89, 90]. As highlighted in Ref. 33, topological edge currents naturally as in Fig. 2(b) reproduce the kinetic ordering of KaiC phosphorylation cycles, although the detailed biophysical mechanisms remained abstract in this work. These missing mechanisms were made concrete (see Fig. 4) in Ref. 40, which also characterized the performance and efficiency of the resulting clock. In particular, it was shown that the key biochemical mechanisms [91, 92, 93, 94] interact via a separation of time-scales to produce the resulting edge currents, and provide the necessary and sufficient conditions for the observed oscillatory cycle [40]. An analysis of the parameter space showed how a variation of the transition rates (e.g. by changing ATP concentration [95] or using a mutant [96]) affects the coherence and dissipation of the oscillation, by tuning the system into and out of the topological transition. This yields key insights into an important regulatory system that had until now required rather complicated models, especially to reproduce the observed kinetic ordering of T and S phosphorylation [97, 98, 99].
On the theoretical level, Ref. 40 characterized the coherence of the resulting cycle, showing that it satisfies theoretical bounds [100] for the most coherent oscillator, equivalent to that of a unicyclic network, without the fine-tuning needed for a unicycle. The topological model shows high coherence compared to other available models, while producing the global day-night cycle with unusually few free parameters. In addition, the coherence and energetic cost of the oscillation were explored using tools from non-equilibrium stochastic thermodynamics, to reveal an efficient regime where coherence increases while cost simultaneously decreases. Lastly, a new indicator of oscillation coherence from spectral gaps in band theory was introduced, to study the saturation of this model on global thermodynamic bounds.
Overall, a topological model is able to explain long-standing puzzles in biology such as how dimensional reduction is achieved in a robust and flexible manner to produce emergent robust oscillations. More generally, it provides design principles for targeted dynamics in synthetic biological systems and in the engineering of reconfigurable materials, e.g., through dissipative self-assembly.
@@ V.1.2 Sensory adaptation
In Ref. 35, the authors used their 1D topological model [Fig. 2(a)] to describe sensory adaption, in particular how E. coli senses chemical concentrations to navigate in response to chemical gradients [101, 102]. Here, the activity of the bacterial sensing complex is modified by its methylation level according to a sigmoid function. More specifically, there is low activity at low methylation, high activity at high methylation, and a narrow region in between where activity changes fast. The sensing complex tends to have a methylation level where activity changes the fastest, which ensures maximal sensitivity to chemoattractant.
@@
V.1.3 Dynamical instability of microtubules
@@
V.2 Synthetic microswimmers
@@ VI.1 Necessity of non-Hermiticity and out-of-equilibriumness
While there is strong motivation to realize topological states in stochastic systems as described above, general principles that govern stochastic topological systems remain lacking, together with an understanding of how stochastic topology generates edge responses necessary for robust function. Recent theoretical work has addressed this question and demonstrated that the only way for stochastic systems to exhibit a localized response is to have a non-Hermitian transition matrix , see Fig. 6 [73]. Note that Hermitian transition matrices always describe equilibrium systems with uniform steady states, whereas non-Hermitian ones can (but not necessarily do) describe out-of-equilibrium ones. This is in line with other phenomena that have no counterparts in equilibrium systems [38, 104, 16].
@@ VI.3 Transfer matrix approach
The key quantities that define a topological state with chiral edge currents in a stochastic systems are the steady state accumulation of probability at the edges of the system, and the steady state edge current, i.e. how fast and in which direction this probability current flows along the edges of the system at steady state, see Fig. 2(b). In Ref. 33, a novel method to obtain these quantities was proposed, based on a transfer matrix approach. The key to the approach is to relate the steady state probabilities of neighboring cells to each other. In particular, for the model in Fig. 2(b), one can obtain a relation , where is a four-component vector containing the probabilities at the four sites of the unit cell that is a distance away from the edge of the system, and is a transfer matrix, independent of . The eigenvalues and eigenvectors of then contain all the necessary information about the edge states. In particular, was found to have two eigenvalues equal to one, corresponding to a uniform steady state of the bulk of the system, and two eigenvalues and , corresponding to edge modes located at opposite edges of the system. From the eigenvector associated to , the probability accumulation and edge current could then be explicitly calculated [33]. Additionally, corresponds to the decay length of the edge mode, which is damped by a factor in a unit cell a distance away from the edge.
@@ VI.4 Topological classification using symmetries
VI.5 Thermodynamic properties and interactions
An interesting alternative approach using thermodynamic properties and the stoichiometric matrix associated to the transition matrix has been found, that can also predict the topological transition [108]. Separately, analysis of the entropy production in a topological model has revealed a regime of high efficiency, in which there is less dissipation despite greater coherence [40].
It would also be potentially useful to consider interactions between several biochemical or ecological systems. Inspired by biological systems under resource constraints, Ref. 33 considered coupled systems with shared boundaries, resulting in (anti)synchronized edge states. Going forward, the effects of different interactions and coupling, which effectively increases the dimension of the system, appears as a promising area for future investigation.
@@ VII Outlook
We have highlighted several open theoretical directions in the previous section. Meanwhile, work remains to be done in the application and verification of existing topological models to realistic settings of biological relevance, as well as in the development of new models that exhibit yet unseen features that are exclusive to stochastic systems and have no quantum counterpart. It would also be useful to analyze the robustness of proposed models to forms of disorder, heterogeneity, and other constraints specified by real-world problems.
@@ Acknowledgements.
- Shankar et al. [2022] S. Shankar, A. Souslov, M. J. Bowick, M. C. Marchetti, and V. Vitelli, Topological active matter, Nature Reviews Physics 2022 4:6 4, 380 (2022).
@@
- Zheng and Tang [2023] C. Zheng and E. Tang, A topological mechanism for robust and efficient global oscillations in biological networks, arXiv:2302.11503 (2023).
@@
- Nelson and Tang [2023] A. Nelson and E. Tang, Non-reciprocity permits edge states and strong localization in stochastic topological systems (2023), arXiv:2310.16720 [cond-mat.stat-mech] .

### v2

@@ III.3 Formalism and mapping to quantum systems
Topological invariants in stochastic systems can be identified using results developed in quantum systems. In Ref. 25, a topological index was identified from the winding number in Fourier space of the phase associated to the determinant of the tilted current matrix associated to the bulk. In follow-up work, this index was further related to the topological index of the Hermitian Hamiltonian of an associated supersymmetric quantum system [26].
A more direct method that flexibly maps many systems described by the master equation to a quantum Hamiltonian was used in Ref. 24, although the resulting Hamiltonian in this case is typically non-Hermitian [71, 72]. Namely, the similarity of the master equation to the Schrödinger equation is exploited, as they are formally identical up to the constant . In quantum systems, topological invariants such as the Chern number are typically calculated from the eigenvectors of the Hamiltonian , see Eq. 2 . When mapping a master equation to a corresponding Schrödinger equation, the eigenvectors of the transition matrix will be identical to those of , in the case of periodic boundary conditions. Hence, the same calculation that verifies if the quantum system is topological using its eigenvectors, can be performed for from a stochastic system. Using the bulk-boundary correspondence, a topological invariant that characterizes the bulk will predict a boundary response for the system. Further works have shown that this bulk-boundary correspondence has additional constraints in stochastic systems as compared to quantum ones [73, 74, 75], which is discussed below in Section VI.
@@ IV Attractive features and useful properties
In the remaining sections of the review, we restrict our focus to stochastic systems, for which there have been a greater number of follow-up works exploring further methods of analysis and potential applications. In particular, while periodic lattices that allow the definition of a bulk and a boundary arise naturally in stochastic systems, the appearance of such lattice structures in ecological networks is less obvious (additionally, realistic population dynamic models do not generally have the conservation laws that allow their reformulation as an ALVE). Nevertheless, many of the conclusions drawn in the following may also be applied to such population dynamics models.
@@ V.1.1 Long timescales: the circadian rhythm
A detailed case study of a topological mechanism in a concrete biological system is that of the KaiABC system for the circadian clock of cyanobacteria [89, 90]. As highlighted in Ref. 24, topological edge currents in Fig. 2(b) naturally reproduce the kinetic ordering of KaiC phosphorylation cycles, although the detailed biophysical mechanisms remained abstract in this work. These missing mechanisms were made concrete (see Fig. 4) in Ref. 38, which also characterized the performance and efficiency of the resulting clock. In particular, it was shown that the key biochemical mechanisms [91, 92, 93, 94] interact via a separation of time-scales to produce the resulting edge currents, and provide the necessary and sufficient conditions for the observed oscillatory cycle [38]. An analysis of the parameter space showed how a variation of the transition rates (e.g. by changing ATP concentration [95] or using a mutant [96]) affects the coherence and dissipation of the oscillation, by tuning the system into and out of the topological transition. This yields key insights into an important regulatory system that had until now required rather complicated models, especially to reproduce the observed kinetic ordering of T and S phosphorylation [97, 98, 99].
On the theoretical level, Refs. 24 and 38 characterized the coherence of the resulting cycle, showing that it satisfies theoretical bounds [100] for the most coherent oscillator, equivalent to that of a unicyclic network, without the fine-tuning needed for a unicycle. The topological model shows high coherence compared to other available models, while producing the global day-night cycle with unusually few free parameters. In addition, the coherence and energetic cost of the oscillation were explored using tools from non-equilibrium stochastic thermodynamics, to reveal an efficient regime where coherence increases while cost simultaneously decreases. This finding provides an experimental prediction where one could test if decreasing ATP concentration increases ADP consumption, which is the opposite of what is usually expected.
Further, cryoEM could be used to distinguish between this topological model and the typical MWC paradigm, since the former predicts new conformational distributions. Specifically, the topological model should show a wide distribution of different A-loop conformational patterns, as opposed to a narrow distribution that is expected from MWC models. One would also expect that mutants mimicking phosphorylated S-sites (T-sites) would correlate with hexamers having mostly buried (exposed) A-loops in the topological model, whereas such correlations would not be expected otherwise. More broadly, a topological model is able to explain long-standing puzzles in biology, such as how dimensional reduction is achieved in a robust and flexible manner to produce emergent robust oscillations.
@@ V.1.2 Sensory adaptation
In Ref. 25, the authors used their 1D topological model [Fig. 2(a)] to describe sensory adaption, in particular how E. coli senses chemical concentrations to navigate in response to chemical gradients [101, 102]. Here, the activity of the bacterial sensing complex is modified by its methylation level according to a sigmoid function. More specifically, there is low activity at low methylation, high activity at high methylation, and a narrow region in between where activity changes fast. The sensing complex tends to have a methylation level where activity changes the fastest, which ensures maximal sensitivity to chemoattractant.
@@
V.1.3 Gene transcription
Topological invariants have also been used to predict different dynamical regimes in a simple gene switch [103] and demonstrate the non-equilibrium phase transitions between these regimes. Here, topology provides new insights where a simple energy landscape paradigm fails: that multistable steady states can lie on the edges of a topological model. In particular, topological analysis reveals a new oscillatory regime in addition to previously recognized bistable and monostable phases [104]. The authors show how local winding numbers predict the steady-state locations in the bistable and monostable phases, while a flux analysis quantifies the respective strengths of steady-state peaks.
This work predicts different transcription regimes (e.g., oscillations vs random switches) that can be distinguished experimentally. Further, it uncovers different mechanisms that lead to multistability or oscillations, that can be useful to quantify existing observations in synthetic gene circuits [105]. Notably, this work develops new methods for systems without a repeated lattice structure or that have disorder, extending the suite of biological networks that can be analyzed using topology [104].
V.1.4 Dynamical instability of microtubules
@@
V.2 Synthetic systems
V.2.1 Microfluidic networks
Due to their tunability and ease of control, microfluidic devices have become natural platforms for the design and realization of different topological states [106]. This includes the patterning of a Lieb lattice in which an active liquid can form a topologically protected sound mode [107], or the construction of chiral topological flows [108].
Active particles can also be guided to accumulate at corners through a second-order skin effect [109], an example of higher-order topological phases, which cause the accumulation of particles at the zero-dimensional corners of a two-dimensional system. More generally, in Ref. 109, a microfluidic device with open chambers connected to each other by narrow connections of various shapes and widths resulted in the active particles jumping stochastically between chambers, thus providing a direct mapping to the systems described in the present review. Microfluidic systems containing active particles therefore offer a versatile testbed for lattices or geometries that support new topological phases and edge states.
V.2.2 Robust microswimmer design
@@ VI.1 Necessity of non-Hermiticity and out-of-equilibriumness
While there is strong motivation to realize topological states in stochastic systems as described above, general principles that govern topological features in stochastic systems, such as the edge responses necessary for robust function, remain lacking. Recent theoretical work has addressed this question and demonstrated that the only way for stochastic systems to exhibit a localized response is to have a non-Hermitian transition matrix , see Fig. 6 [73]. Note that Hermitian transition matrices always describe equilibrium systems with uniform steady states, whereas non-Hermitian ones can (but not necessarily do) describe out-of-equilibrium ones. This is in line with other phenomena that have no counterparts in equilibrium systems [56, 111, 37].
@@ VI.3 Transfer matrix approach
The key quantities that define a topological state with chiral edge currents in a stochastic systems are the steady state accumulation of probability at the edges of the system, and the steady state edge current, i.e. how fast and in which direction this probability current flows along the edges of the system at steady state, see Fig. 2(b). In Ref. 24, a novel method to obtain these quantities was proposed, based on a transfer matrix approach. The key to the approach is to relate the steady state probabilities of neighboring cells to each other. In particular, for the model in Fig. 2(b), one can obtain a relation , where is a four-component vector containing the probabilities at the four sites of the unit cell that is a distance away from the edge of the system, and is a transfer matrix, independent of . The eigenvalues and eigenvectors of then contain all the necessary information about the edge states. In particular, was found to have two eigenvalues equal to one, corresponding to a uniform steady state of the bulk of the system, and two eigenvalues and , corresponding to edge modes located at opposite edges of the system. From the eigenvector associated to , the probability accumulation and edge current could then be explicitly calculated [24]. Additionally, corresponds to the decay length of the edge mode, which is damped by a factor in a unit cell a distance away from the edge.
@@ VI.4 Topological classification using symmetries
VI.5 Stochastic thermodynamics
An interesting alternative approach using thermodynamic properties and the stoichiometric matrix associated to the transition matrix has been suggested, that can also predict the topological transition [115]. Separately, analysis of the entropy production in a topological model has revealed a regime of high efficiency, in which there is less dissipation despite greater coherence [38]. Stochastic thermodynamics and the associated fluctuation theorems [116] can also be used to infer the non-equilibrium driving forces (entropy production) from observable trajectories, sometimes with close connection to the underlying topology [117].
VI.6 Dynamical phases in interacting systems
Besides the population dynamics models described in Section III.2, which corresponded to the mean field level of an interacting system, most of the works discussed so far concerned noninteracting systems, i.e. effectively single-particle models in one- or two-dimensional lattices. One particularly interesting direction for future work is the application of topological tools to interacting, many-particle stochastic systems. Inspired by biological systems under resource constraints, Ref. 24 considered two coupled systems with shared boundaries, resulting in (anti)synchronized edge states that were studied numerically.
Another recent work [118] explored a paradigmatic example of interacting dynamics, namely the symmetric simple exclusion process, and showed how symmetry-protected topological phases can be used to identify distinct dynamical phases with regards to the types of rare trajectories that emerge in the system. Future work in this direction could explore genuinely non-equilibrium interacting processes, such as asymmetric exclusion processes, and apply the same tools to interacting generalizations of systems in which even the single particle dynamics is topological, such as those described above.
@@ VII Outlook
While we have highlighted possible experimental realizations in synthetic gene circuits or single-molecule measurements, future work remains to be done in the verification of existing topological models for settings of biological relevance. On the theoretical side, we reviewed several developments and advances in the previous section. Still, we look forward to the development of new models that could exhibit as yet unknown features that are exclusive to stochastic systems and have no quantum counterpart. Lastly, while there has been initial progress in developing tools amenable to systems with disorder or heterogeneity [104], more work needs to be done to analyze the robustness of existing models to other forms of disorder or constraints imposed by real-world scenarios.
It should be noted that most of the works described here dealt with one- or two-dimensional systems. Most biomolecular and ecological phenomena, on the other hand, take place in much higher-dimensional spaces. Future work should therefore explore topological phenomena in high-dimensional systems and, relatedly, the higher-order topological phenomena that will be needed to localize the system’s dynamics to much lower dimensional manifolds. While significant progress has been made in understanding higher-order topological phases in quantum systems [119, 120], very little is known about their emergence in classical stochastic systems.
@@ Acknowledgements.
- Shankar et al. [2022a] S. Shankar, A. Souslov, M. J. Bowick, M. C. Marchetti, and V. Vitelli, Topological active matter, Nature Reviews Physics 2022 4:6 4, 380 (2022a).
- Zheng and Tang [2024] C. Zheng and E. Tang, A topological mechanism for robust and efficient global oscillations in biological networks, Nature Communications 2024 15:1 15, 1 (2024).
@@
- Nelson and Tang [2024] A. Nelson and E. Tang, Nonreciprocity is necessary for robust dimensional reduction and strong responses in stochastic topological systems, Physical Review B 110, 155116 (2024).
@@
- Schultz et al. [2007] D. Schultz, J. N. Onuchic, and P. G. Wolynes, Understanding stochastic simulations of the smallest genetic networks, The Journal of Chemical Physics 126, 245102 (2007).
- Nelson et al. [2025] A. Nelson, P. Wolynes, and E. Tang, Topology of the simplest gene switch, arXiv preprint arXiv:2503.01717 (2025).
- Zhu et al. [2022] R. Zhu, J. M. Del Rio-Salgado, J. Garcia-Ojalvo, and M. B. Elowitz, Synthetic multistability in mammalian cells, Science 375, eabg9765 (2022).
- Shankar et al. [2022b] S. Shankar, A. Souslov, M. J. Bowick, M. C. Marchetti, and V. Vitelli, Topological active matter, Nature Reviews Physics 4, 380 (2022b).
- Souslov et al. [2017] A. Souslov, B. C. Van Zuiden, D. Bartolo, and V. Vitelli, Topological sound in active-liquid metamaterials, Nature Physics 2017 13:11 13, 1091 (2017).
- Copar, Simon and Kos, Ziga and Emersic, Tadej and Tkalec, Uros [2020] Copar, Simon and Kos, Ziga and Emersic, Tadej and Tkalec, Uros, Microfluidic control over topological states in channel-confined nematic flows, Nature Communications 2020 11:1 11, 1 (2020).
- Palacios et al. [2021] L. S. Palacios, S. Tchoumakov, M. Guix, I. Pagonabarraga, S. Sánchez, and A. G. Grushin, Guided accumulation of active particles by topological design of a second-order skin effect, Nature Communications 12, 1 (2021).
@@
- Seifert [2012] U. Seifert, Stochastic thermodynamics, fluctuation theorems and molecular machines, Reports on progress in physics 75, 126001 (2012).
- Mahault et al. [2022] B. Mahault, E. Tang, and R. Golestanian, A topological fluctuation theorem, Nature Communications 13, 3036 (2022).
- Garrahan and Pollmann [2024] J. P. Garrahan and F. Pollmann, Topological phases in the dynamics of the simple exclusion process, Phys. Rev. E 109, L032105 (2024).
- Xie et al. [2021] B. Xie, H.-X. Wang, X. Zhang, P. Zhan, J.-H. Jiang, M. Lu, and Y. Chen, Higher-order band topology, Nature Reviews Physics 3, 520 (2021).
- Yang et al. [2024] Y.-B. Yang, J.-H. Wang, K. Li, and Y. Xu, Higher-order topological phases in crystalline and non-crystalline systems: a review, Journal of Physics: Condensed Matter 36, 283002 (2024).